OpenStudy (anonymous):

solve sec^2(x) over the interval [-pi/2, 5pi/2] I am stumped with this one at the moment - thanks for any help!

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

solve sec^2(x)..wts that mean?

OpenStudy (toxicsugar22):

hi

OpenStudy (anonymous):

do you want us to integrate?

OpenStudy (anonymous):

or solve sec^2(x)= (blank)

OpenStudy (anonymous):

not sure: here is one of the multiple choices that I chose (as wrong) {0,5/2pi}

OpenStudy (anonymous):

oh my bad; sec^2(x)=1

OpenStudy (anonymous):

ok that makes more sense.

OpenStudy (anonymous):

haha! sorry

OpenStudy (anonymous):

I am really rusty with my unit circle trig if you can explain how you do it too that would be a huge help!

OpenStudy (anonymous):

sec^2(x)=1/(cos^2(x))=1 cos^2(x)=1 cos(x)=+/- 1 x=Pi *n where n is an integer. so find the multiples of Pi that are in that interval

OpenStudy (anonymous):

{0, Pi, 2Pi}

OpenStudy (anonymous):

I got to the part just before cos(x)=+/- 1; how can you change cos^2(x)=1 like you did? thanks

OpenStudy (anonymous):

square root of both sides

OpenStudy (anonymous):

haha - thanks!

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

so you simply reduced it to the point that you could plot it on the unit circle?

OpenStudy (anonymous):

it basicaly means cos(x)=+1 or -! solution. so cosx takes those values for 0 ,pi ,2pi

OpenStudy (anonymous):

i reduced it to a trig equation I know the result of. cos(x) takes on 1 and -1 at any multiple of Pi

OpenStudy (anonymous):

cosx=1/secx and so it is posible to write cosx=+/-1 from sec^2(x)=+/-1

OpenStudy (anonymous):

Thank, that all makes sense: how did you interpret the range part of the question though?

OpenStudy (anonymous):

i found all the integer multiples of Pi that fall within the given interval

OpenStudy (anonymous):

3Pi was too large, -1 Pi too small, 0, Pi, 2Pi work

OpenStudy (anonymous):

gotcHA

OpenStudy (anonymous):

out of interest how would 3 pi plot on the unit circle - if there was no range given in the question?

OpenStudy (anonymous):

would the answer change if the question were sin(x)+/- Pi?

OpenStudy (anonymous):

Thank you all for your help!